# Isotropic Covariance Functions for Disparate Spatial Resolutions

### v2

@@ 4 Numerical Examples
where $h$ represents the distance between two elements from the index set, $\phi$ controls the strength of the spatial dependence, and $\nu$ governs smoothness. Both families belong to the class $\Phi_{\infty}$ and are therefore valid on Hilbert spaces. For interpretability, dependence is summarized using the practical range $\rho$, defined as the distance where correlation decays to 0.1. For the PEXP family, $\rho={\log(10)}^{1/\nu}\phi$ (Godoy et al.,, 2025), while for the Matérn family, $\rho\approx\phi\sqrt{8\nu}$ (Lindgren et al.,, 2011). These models form the basis for distance comparisons and eigenvalue-based diagnostics that follow.
@@
We computed pairwise distance matrices for all configurations shown in Figure 1 using both the ball-Hausdorff and Hausdorff distances. Distances were rescaled to the unit interval by dividing by the maximum observed distance within each scenario, so that comparisons reflect relative rather than absolute scale differences across spatial configurations. Under this joint normalization, the values for of either metric need not span the full unit interval, as the observed minimum and maximum can be bounded away from 0 and 1 by geometric constraints inherent to the synthetic designs, such as shared boundaries, partial overlap, and limited variability in set geometry. Figure 2 displays two-dimensional kernel density estimates comparing the resulting distance values. The concentration of mass along the $y=x$ diagonal indicates close empirical agreement between the two distances, with no systematic curvature or dispersion as spatial complexity increases from gridded to fused and irregular areal units. In particular, the near-monotone relationship across all configurations suggests that the ball-Hausdorff distance preserves the relative proximity structure induced by the Hausdorff distance, despite relying on a substantially reduced geometric representation.
@@
We conclude the numerical examples with an eigenvalue-based comparison of covariance matrices constructed using the ball-Hausdorff and standard Hausdorff distances. Two covariance families are considered: the PEXP family with $\nu=1$ and the Matérn family with $\nu=2.5$. For the Matérn family, covariance matrices are constructed using square-root transformed distances, as required by Theorem 2.3. The practical range parameter is varied over $\rho\in[0.8,1.5]$, where the difference between the positive definiteness of the resulting covariance matrices from the two distances are notable. The smallest eigenvalue of each resulting covariance matrix, denoted $\lambda_{1}$, is used to assess positive definiteness, with negative values indicating invalid covariance structures.
@@
Figure 3 summarizes the resulting eigenvalue diagnostics across spatial configurations and domains. In all cases considered, covariance matrices based on the ball-Hausdorff distance remain positive definite, with $\lambda_{1}\geq 0$ throughout the examined range of $\rho$. In contrast, covariance matrices constructed using the standard Hausdorff distance frequently exhibit negative eigenvalues, even when square-root transformation is applied for the Matérn family, a reported limitations of Hausdorff-based constructions in set-indexed settings (Godoy et al.,, 2025). These violations are most pronounced for fused data and for configurations defined on the sphere. Collectively, these results serve as a final numerical confirmation that replacing the Hausdorff distance with the ball-Hausdorff distance leads to systematically improved validity of covariance constructions in set-indexed spatial settings.
@@ 5 Discussion
The application of covariance families under the proposed framework depends on the scale at which the ball–Hausdorff distance is transformed to preserve positive definiteness. While the PEXP family with $\nu\in(0,1]$ can be applied directly to the ball-Hausdorff distance, other families, such as the Matérn class, require the square-root transformed distance to ensure theoretical validity (Theorem 2.3). Because the square-root is a monotone transformation, inference on range-related parameters remains identifiable on the original distance scale. In frequentist settings, invariance of the maximum likelihood estimator permits reporting the practical range on the original scale, with uncertainty quantified, for example, via the delta method. In Bayesian analyses, the same transformation may be applied to posterior samples or handled explicitly through a Jacobian adjustment. These properties allow direct construction of valid set-indexed Gaussian processes for problems such as spatial data fusion and change-of-support, while still accommodating stochastic integrals when scientifically appropriate. An interesting theoretical direction is the development of valid cross-covariance functions for sets (Genton and Kleiber,, 2015), which would support modeling when responses and explanatory variables are observed on different spatial supports.
Several theoretical questions remain open. One concerns the mean-square differentiability of random fields defined through the proposed covariance construction. For the standard Matérn family, differentiability is governed by the smoothness parameter (Stein,, 1999, pg. 32), but extending this notion to set-indexed or manifold-indexed processes requires more general concepts of differentiation (Khastan et al.,, 2021). This issue is non-trivial, even for Gaussian processes on spheres (Alegría et al.,, 2021; Guinness and Fuentes,, 2016), and while recent work has defined mean-square differentiability for Gaussian processes on manifolds using smooth paths (Lin et al.,, 2019), a systematic treatment in the present setting remains to be developed. A further limitation arises from the pseudometric nature of the ball–Hausdorff distance: distinct sets with identical centers and radii are indistinguishable, which may lead to singular covariance matrices in degenerate cases. As in many reduced-representation models, this issue can be addressed by introducing a nugget effect, a standard regularization in geostatistical modeling that accounts for measurement error and unresolved small-scale variability (Cressie,, 1993, pg. 59–60).
@@ Acknowledgments
Appendix A Proofs
A.1 Lemma 2.1
Proof.
$d(x,u)\leq d(x,v)+d(v,u)<k+r,$
thus, $u\in\mathrm{B}_{r+k}(x)$.
A.2 Corollary 2.2.1
Proof.
This establishes that $\exp\{-\kappa d^{\nu}\}$ is RPD on $(D,d)$. ∎
Appendix B Length and geodesic spaces
@@
Definition B.1 (Definition 2.3.1 in Burago et al., (2001)).
@@
When the construction above induces the initial metric $d$ (i.e., $d_{L}=d$), than it is said to be intrinsic (Burago et al.,, 2001).
@@
Appendix C An upper bound for the ball-Hausdorff distance

### v3

@@ 4 Numerical Examples
where $h$ represents the distance between two elements from the index set, $\phi$ controls the strength of the spatial dependence, and $\nu$ governs smoothness. Both families belong to the class $\Phi_{\infty}$ and are therefore valid on Hilbert spaces. For interpretability, dependence is summarized using the practical range $\rho$, defined as the distance where correlation decays to 0.1. For the PEXP family, $\rho={\log(10)}^{1/\nu}\phi$ (Godoy et al., 2026a, ), while for the Matérn family, $\rho\approx\phi\sqrt{8\nu}$ (Lindgren et al.,, 2011). These models form the basis for distance comparisons and eigenvalue-based diagnostics that follow.
@@
We computed pairwise distance matrices for all configurations shown in Figure 1 using both the ball-Hausdorff and Hausdorff distances. Distances were rescaled to the unit interval by dividing by the maximum observed distance within each scenario, so that comparisons reflect relative rather than absolute scale differences across spatial configurations. Figure 2 displays two-dimensional kernel density estimates comparing the resulting distance values. The concentration of mass along the $y=x$ diagonal indicates close empirical agreement between the two distances, with no systematic curvature or dispersion as spatial complexity increases from gridded to fused and irregular areal units. In particular, the near-monotone relationship across all configurations suggests that the ball-Hausdorff distance preserves the relative proximity structure induced by the Hausdorff distance, despite relying on a substantially reduced geometric representation.
@@
We conclude the numerical examples with an eigenvalue-based comparison of covariance matrices constructed using the ball-Hausdorff and standard Hausdorff distances. Two covariance families are considered: the PEXP family with $\nu=1$ and the Matérn family with $\nu=2.5$. For the Matérn family, covariance matrices are constructed using square-root transformed distances, as required by Theorem 2.3. The practical range parameter is varied over $\rho\in[0.8,1.5]$, where the differences between the positive definiteness of the resulting covariance matrices from the two distances are notable. The smallest eigenvalue of each resulting covariance matrix, denoted $\lambda_{1}$, is used to assess lack of positive definiteness, with negative values indicating invalid covariance structures.
@@
Figure 3 summarizes the resulting eigenvalue diagnostics across spatial configurations and domains. In all cases considered, covariance matrices based on the ball-Hausdorff distance remain positive definite, with $\lambda_{1}\geq 0$ throughout the examined range of $\rho$. In contrast, covariance matrices constructed using the standard Hausdorff distance frequently exhibit negative eigenvalues, even when square-root transformation is applied for the Matérn family, a reported limitation of Hausdorff-based constructions in set-indexed settings (Godoy et al., 2026a, ). These violations are most pronounced for fused data and for configurations defined on the sphere. Collectively, these results serve as a final numerical confirmation that replacing the Hausdorff distance with the ball-Hausdorff distance leads to systematically improved validity of covariance constructions in set-indexed spatial settings.
@@ 5 Discussion
The application of covariance families under the proposed framework depends on the scale at which the ball–Hausdorff distance is transformed to preserve positive definiteness. While the PEXP family with $\nu\in(0,1]$ can be applied directly to the ball-Hausdorff distance, other families, such as the Matérn class, require the square-root transformed distance to ensure theoretical validity (Theorem 2.3). Because the square-root is a monotone transformation, inference on range-related parameters remains identifiable on the original distance scale. In frequentist settings, invariance of the maximum likelihood estimator permits reporting the practical range on the original scale, with uncertainty quantified, for example, via the delta method. In Bayesian analyses, the same transformation may be applied to posterior samples or handled explicitly through a Jacobian adjustment. These properties allow direct construction of valid set-indexed Gaussian processes for problems such as spatial data fusion and change-of-support, while still accommodating stochastic integrals when scientifically appropriate.
Several theoretical questions remain open. One concerns the mean-square differentiability of random fields defined through the proposed covariance construction. For the standard Matérn family, differentiability is governed by the smoothness parameter (Stein,, 1999, pg. 32), but extending this notion to set-indexed or manifold-indexed processes requires more general concepts of differentiation (Khastan et al.,, 2021). This issue is non-trivial, even for Gaussian processes on spheres (Alegría et al.,, 2021), and while recent work has defined mean-square differentiability for Gaussian processes on manifolds using smooth paths (Lin et al.,, 2019), a systematic treatment in the present setting remains to be developed. A further limitation arises from the pseudometric nature of the ball–Hausdorff distance: distinct sets with identical centers and radii are indistinguishable, which may lead to singular covariance matrices in degenerate cases. As in many reduced-representation models, this issue can be addressed by introducing a nugget effect, a standard regularization in geostatistical modeling that accounts for measurement error and unresolved small-scale variability (Cressie,, 1993, pg. 59–60).
@@ Acknowledgments
Appendix A Length and geodesic spaces
@@
Definition A.1 (Definition 2.3.1 in Burago et al., (2001)).
@@
When the construction above induces the initial metric $d$ (i.e., $d_{L}=d$), then it is said to be intrinsic (Burago et al.,, 2001).
@@
Appendix B An upper bound for the ball-Hausdorff distance